# Toward Explainable Offline RL: Analyzing Representations in Intrinsically Motivated Decision Transformers

### v1

@@ 4.3 Performance Results
Table 1 presents performance results across both medium and medium-replay datasets, where intrinsic motivation variants demonstrate environment-specific effectiveness patterns.
On medium datasets, EDT-TIL achieved the best performance in 2 out of 4 environments (Walker2d: 73.50 vs 68.50 HNS; Hopper: 59.63 vs 57.49/59.31 HNS for baseline/SIL).
@@ 4.4 Embedding Analysis Results
Table 2 reveals the mechanistic basis for performance improvements through analysis of embedding properties. Each environment exhibits a distinct correlation pattern between representational metrics and performance. More specifically we can see the following:
  • Ant: this environment shows a strong negative correlation with covariance trace (-0.907). This might suggest that reduced total variance distribution improves performance.
  • HalfCheetah: in this case we can see a positive correlation with covariance trace (0.850), suggesting that increased representational capacity benefits this environment.
  • Hopper: A positive correlation with cosine similarity (+0.658), suggests that increased similarity between state representations enhances performance
  • Walker2d: In this case a strong negative correlation with cosine similarity (-0.950), indicating that increased orthogonality between embeddings is crucial.
Such environment-specific patterns demonstrate that intrinsic motivation mechanisms create tailored representational structures aligned with task demands and consistent with the biological principle of adaptive representational organization, which we hypothesized in our biological plausible framework.
Examining the embedding properties across models, we can further see and interesting effect. More specifically the analysis reveals distinct representational effects of each intrinsic motivation variant. EDT-SIL consistently creates more compact representations through reduced covariance trace and L2 norms. EDT-TIL promotes representational orthogonality via reduced cosine similarity (Walker2d: -0.950; Hopper: +0.658), demonstrating environment-specific optimization strategies that mirror biological neural decorrelation principles. The complementary nature of those mechanisms suggests that different intrinsic motivation approaches implement distinct aspects of biological representational regulation. EDT-SIL enhances representational efficiency at the input level, while EDT-TIL optimizes sequential processing structures. This division of regulatory functions aligns with hierarchical organization principles observed in biological neural systems, where different processing stages maintain distinct homeostatic mechanisms.
These findings strongly suggests that intrinsic motivation in EDTs operates as more than a simple exploration bonus: it acts as a representational prior that shapes embedding geometry in biologically plausible ways, creating environment-specific organizational structures that facilitate improved decision-making.
@@ 5 Conclusions
In our work we investigated how, introducing intrinsic motivation mechanisms in Elastic Decision Transformers, could shape learned representations and their correlation with task performance across multiple continuous control environments. Our systematic analysis revealed several key findings, that could suggest the relationship between empirical performance improvements and the underlying biologically inspired mechanisms. In a first instance, our experiments demonstrate that intrinsic motivation variants (EDT-SIL and EDT-TIL) consistently outperform the baseline EDT across most environments and datasets, with notable improvements on medium datasets. The 3-layers RND configuration emerged as optimal, aligning with biological principles of representational balance where intermediate capacity avoids both under-representation and overfitting. Moreover, the post-hoc explainability analysis reveals that each environment exhibits distinct patterns in how embedding properties correlate with performance. EDT-SIL creates compact representations (reduced covariance/L2 norms), while EDT-TIL shows environment-specific patterns (Walker2d: orthogonality; Hopper: inter-dimensional correlations). EDT-TIL promotes representational orthogonality through reduced cosine similarity, particularly evident in Walker2d where this property shows the strongest correlation with performance (-0.950).
@@
Such findings show that intrinsic motivation operates as more than a simple exploration bonus: it acts as a representational prior that shapes embedding geometry in biologically plausible ways. The complementary nature of EDT-SIL and EDT-TIL mechanisms mirrors the hierarchical organization principles observed in biological neural systems, where different processing stages maintain distinct homeostatic mechanisms for representational regulation. The environment-specific correlation patterns between embedding metrics and performance suggest that intrinsic motivation mechanisms create tailored representational structures aligned with task demands. This adaptive representational organization provides a mechanistic explanation for why intrinsic motivation improves performance beyond simple reward optimization, supporting the biological principle that successful adaptation emerges from appropriate representational regulation. Limitations include a small number of environment tested. In future work we could extend the framework to analyze transformer outputs and action representations, providing a more comprehensive view of how intrinsic motivation affects different components of the decision-making process. Additionally, analyzing the temporal dynamics of embedding evolution during the training process, could provide deeper insights into how intrinsic motivation shapes learning trajectories and representational development over time.
Appendix A Extended Embedding Analysis
A.1 Detailed Medium-Replay Dataset Analysis
@@
Appendix B Implementation Details
B.1 Training Configuration
@@
The 3-layer RND configuration consists of a predictor network (Linear(input_size, 512) $\rightarrow$ ELU $\rightarrow$ Linear(512, 512) $\rightarrow$ ELU $\rightarrow$ Linear(512, 512)) and a target network (Linear(input_size, 512)). All linear layers use orthogonal initialization with gain $\sqrt{2}$ and zero bias initialization as in [5]. The target network remains frozen during training. Training requires approximately 1 hour and 3GB VRAM on an NVIDIA RTX 2080Ti system with Intel i9-9820X processor and 64GB RAM. EDT-SIL and EDT-TIL variants introduce negligible computational overhead compared to baseline EDT.
@@
Appendix C Statistical Analysis Details
C.1 ANOVA Results
Environment Metric F-statistic $p$-value
Medium Dataset
Ant cov_trace 10.69 0.011
l2_norm_mean 11.25 0.009
HalfCheetah cov_trace 380.16 $<$0.001
cosine_similarity_mean 701.46 $<$0.001
Hopper cov_trace 351.71 $<$0.001
l2_norm_mean 325.74 $<$0.001
Walker2d cosine_similarity_mean 56.16 $<$0.001
Medium-Replay Dataset
HalfCheetah cov_trace 973.83 $<$0.001
Hopper l2_norm_mean 53.87 $<$0.001
Walker2d cosine_similarity_mean 15.34 0.004

### v2

@@ 3.1 Performance Results
Table 1 presents performance results across both medium and medium-replay datasets, where intrinsic motivation variants demonstrate environment-specific effectiveness patterns. On medium datasets, EDT-TIL achieved the best performance in 2 out of 4 environments (Walker2d: 73.50 vs 68.50 HNS; Hopper: 59.63 vs 57.49/59.31 HNS for baseline/SIL).
@@ 3.2 Embedding Analysis Results
Table 2 reveals the mechanistic basis for performance improvements through analysis of embedding properties. Each environment exhibits a distinct correlation pattern between representational metrics and performance. Ant shows a strong negative correlation with covariance trace (-0.907), suggesting that reduced total variance distribution improves performance. HalfCheetah exhibits a positive correlation with covariance trace (0.850), indicating that increased representational capacity benefits this environment. Hopper demonstrates a positive correlation with cosine similarity (+0.658), suggesting that increased similarity between state representations enhances performance. Walker2d shows a strong negative correlation with cosine similarity (-0.950), indicating that increased orthogonality between embeddings is crucial. Such environment-specific patterns demonstrate that intrinsic motivation mechanisms create tailored representational structures aligned with task demands and consistent with the biological principle of adaptive representational organization.
Examining the embedding properties across models, we can further see and interesting effect. More specifically the analysis reveals distinct representational effects of each intrinsic motivation variant. EDT-SIL consistently creates more compact representations through reduced covariance trace and L2 norms. EDT-TIL promotes representational orthogonality via reduced cosine similarity (Walker2d: -0.950; Hopper: +0.658), demonstrating environment-specific optimization strategies that mirror biological neural decorrelation principles. The complementary nature of those mechanisms suggests that different intrinsic motivation approaches implement distinct aspects of biological representational regulation. EDT-SIL enhances representational efficiency at the input level, while EDT-TIL optimizes sequential processing structures. This division of regulatory functions aligns with hierarchical organization principles observed in biological neural systems, where different processing stages maintain distinct homeostatic mechanisms. These findings may suggest that intrinsic motivation may help in acting as a representational prior, shaping the embedding geometry.
@@ 4 Conclusions
In our work we investigated how introducing intrinsic motivation mechanisms in Elastic Decision Transformers shapes learned representations and their correlation with task performance across multiple continuous control environments. Our systematic analysis revealed several key findings that suggest the relationship between empirical performance improvements and underlying mechanisms. Our experiments demonstrate that intrinsic motivation variants (EDT-SIL and EDT-TIL) consistently outperform the baseline EDT across most environments and datasets, with the 3-layer RND configuration emerging as optimal. The post-hoc explainability analysis reveals that each environment exhibits distinct patterns in how embedding properties correlate with performance. EDT-SIL creates compact representations through reduced covariance and L2 norms, while EDT-TIL promotes representational orthogonality through reduced cosine similarity, particularly evident in Walker2d. These findings show that intrinsic motivation might operate as more than a simple exploration bonus: it acts as a representational prior that shapes embedding geometry in biologically plausible ways. The complementary nature of EDT-SIL and EDT-TIL mechanisms mirrors hierarchical organization principles observed in biological neural systems, where different processing stages might maintain distinct homeostatic mechanisms. The environment-specific correlation patterns suggest that intrinsic motivation mechanisms create tailored representational structures aligned with task demands, providing a mechanistic explanation for why intrinsic motivation improves performance beyond simple reward optimization. Regarding the limitations of our study, we primarily focused on the geometrical properties of the embeddings rather than on explicit explainability measures. Further work could include application of explainability models to understand the impact of certain features and vector dimensionality reduction on the final performances of the models proposed. Moreover the initial variations in HNS show very small variations (of the order of percent). This is of course an initial assessment, which should be further exploited and tested, to further prove the significance of the results obtained. In addition, future work could extend the framework to analyze transformer outputs and action representations, and investigate temporal dynamics of embedding evolution during training to understand how intrinsic motivation shapes learning trajectories over time.
Appendix A Biological Plausibility
In this section we propose some insights on the biological plausibility of the proposed model.
A.1 Homeostatic Regulation: towards a biologically inspired reinforcement learning approach
The quest for explainable AI in reinforcement learning (RL) has a tight parallel with biological learning, where intrinsic motivation shapes adaptive behavior through allostatic regulation, achieving stability by adjusting predictions rather than fixing a–priori parameters [allostasi1988]. Such predictive adaptation is reflected in how the so called Random Network Distillation (RND) models operate: learning is driven by discrepancies between predicted and actual inputs, echoing brain mechanisms that constantly update internal models based on sensory prediction errors [friston2010free, Clark_2013]. These prediction hierarchies span from primary sensory to higher-order cortical processing [friston2009predictive], suggesting intrinsic motivation can be applied across representational levels. Moreover, biological systems maintain representational homeostasis, optimizing information processing through the regulation of capacity and structural organization [zenke2017, ganguli2012compressed]. Intrinsic motivation fosters flexible learning and generalization [oudeyer2007intrinsic]. In transformer-based models like Elastic Decision Transformers, auxiliary losses based on RND act as allostatic regulators, guiding representational structure without altering offline reward signals. Such losses can help preventing representational collapse, mirroring biological mechanisms that sustain learning adaptability and predictive efficiency.
Appendix B Random Network Distillation
In this section, we detail how the RND module is integrated into the EDT architecture. We further present the analysis conducted to find the optimal number of layers in RND networks.
B.1 RND Architecture Analysis
Figure 1 shows both EDT-SIL and EDT-TIL variants, where the RND module operates on state embeddings or transformer outputs respectively. The dashed lines indicate backpropagation paths for each variant, with $L_{int}$ from the RND module contributing to the total loss $L_{EDT}$ alongside standard prediction losses. The target network of the RND module has frozen weights and is never updated, as proposed in [rnd].
Figure 1: Architecture of the Elastic Decision Transformer with intrinsic motivation mechanisms proposed in [esann2025].
B.2 RND Layer Configuration Analysis
To optimize the intrinsic motivation mechanism and understand the impact of RND network capacity on intrinsic motivation effectiveness, we conducted a systematic investigation of the predictor network depth on both performance and representation quality. The motivation for this analysis stems from the hypothesis that different network capacities may capture different levels of representational complexity, potentially affecting both the quality of intrinsic rewards and the resulting policy performance. We evaluated three RND predictor configurations: a 1-layer RND (considering it as a minimal architecture to establish a baseline for intrinsic reward generation), a 3-layers RND as the default configuration in [esann2025], and a 10-layers RND, considering it as a high-capacity variant to test whether increased expressiveness improves intrinsic motivation. This analysis was conducted exclusively on our best performing dataset (i.e. Medium Datasets), as these demonstrated the most promising initial results, in order to optimize the RND predictor network architecture.
The 3-layer configuration emerged as optimal across both EDT-SIL and EDT-TIL variants. Figure 2 demonstrates that 3-layer variants (highlighted with red borders) consistently achieve the highest cumulative scores across all environments. This finding aligns with the biological principle of representational balance: too few layers (1-layer) may lack sufficient capacity to capture complex predictive relationships, while too many layers (10-layer) may lead to overfitting or representational instability.
Appendix C Extended Embedding Analysis
In this section, we detail the metrics used to analyze embeddings and provide an in-depth analysis of the properties of embeddings in medium-replay datasets.
C.1 Embedding Characterization Framework
  • Covariance Trace: This metric measures the total variance distributed across embedding dimensions as per:
    $cov\_trace=\mathrm{Tr}(\mathrm{Cov}(E))$ (2)
  • L2 Norm: The mean magnitude of embedding vectors quantifies representational compactness, as per:
    $l2\_norm=\frac{1}{N}\sum_{i=1}^{N}|e_{i}|_{2}$ (3)
  • Cosine Similarity: The average pairwise cosine similarity within each tensor assesses representational orthogonality as per:
C.2 Detailed Medium-Replay Dataset Analysis
@@
Appendix D Statistical Analysis Details
D.1 ANOVA Results
Environment Metric F-statistic $p$-value
Medium Dataset
Ant cov_trace 10.69 0.011
l2_norm_mean 11.25 0.009
HalfCheetah cov_trace 380.16 <<0.001
cosine_similarity_mean 701.46 <<0.001
Hopper cov_trace 351.71 <<0.001
l2_norm_mean 325.74 <<0.001
Walker2d cosine_similarity_mean 56.16 <<0.001
Medium-Replay Dataset
HalfCheetah cov_trace 973.83 <<0.001
Hopper l2_norm_mean 53.87 <<0.001
Walker2d cosine_similarity_mean 15.34 0.004
Appendix E Implementation Details
E.1 Training Configuration
@@
The 3-layer RND configuration consists of a predictor network (Linear(input_size, 512) $\rightarrow$ ELU $\rightarrow$ Linear(512, 512) $\rightarrow$ ELU $\rightarrow$ Linear(512, 512)) and a target network (Linear(input_size, 512)). All linear layers use orthogonal initialization with gain $\sqrt{2}$ and zero bias initialization as in [rnd]. The target network remains frozen during training. Training requires approximately 1 hour and 3GB VRAM on an NVIDIA RTX 2080Ti system with Intel i9-9820X processor and 64GB RAM. EDT-SIL and EDT-TIL variants introduce negligible computational overhead compared to baseline EDT.